LUCIFER Commissioning - 2008 December 13 UT

Observers: W. Seifert, N. Ageorges, M. Jutte, J. Hill (on mtn)
Telescope Operator: W. Wack
Software Support: P. Grenz (on call)


Night lost to clouds.



02:10 Making tests presets with dome closed to exercise new GCS and PCS builds

02:30 Tracking with dome closed

04:00 Had to force Ta 11.70 (17) to get pressure up to 126 Bar to avoid getting stuck at EL=75 deg.


04:10 Open dome under heavy cirrus clouds. We do see bright stars, and we have some quick tests to verify offsetting and guiding. T=-0 degC D=-6 degC

04:16 See an actual star on LUCIFER IE=+9.1 CA=+4.3

04:18 preset ACTIVE, but GCS couldn't see the star. Clouds.

04:23 preset ACTIVE to 7.7 mag (r) star

04:26 preset ACTIVE to 0 mag star. Still too many clouds.


04:32 - Close dome. It was fun while it lasted.

05:45 - Snow is beginning to fall.

-- JohnHill - 13 Dec 2008
Topic revision: r2 - 13 Dec 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback